### v1

# Relative Camera Pose Regression

Abstract
Pre-training techniques have greatly advanced computer vision, with CroCo’s cross-view completion approach yielding impressive results in tasks like 3D reconstruction and pose regression. However, this method requires substantial overlap between training pairs, limiting its effectiveness. We introduce Alligat0R, a novel pre-training approach that reformulates cross-view learning as a co-visibility segmentation task. Our method predicts whether each pixel in one image is co-visible in the second image, occluded, or outside the field of view (FOV), enabling the use of image pairs with any degree of overlap and providing interpretable predictions. To support this, we present Cub3, a large-scale dataset with 2.5 million image pairs and dense co-visibility annotations derived from the nuScenes dataset. This dataset includes diverse scenarios with varying degrees of overlap. The experiments show that Alligat0R significantly outperforms CroCo in relative pose regression, especially in scenarios with limited overlap. Alligat0R and Cub3 will be made publicly available.
Figure 1: We introduce Alligat0R, a pretraining method for binocular vision robust to wide baselines. Alligat0R explicitly segments pixels as co-visible, occluded, or outside field-of-view, overcoming the fundamental limitation of CroCo [45, 46] which attempts to reconstruct potentially unobservable regions. We demonstrate Alligat0R on relative pose regression and show that it allows us to significantly outperform the state-of-the-art, especially on pairs with very low overlap.
1 Introduction
Pre-training techniques have revolutionized computer vision by enabling large models to learn rich representations [3, 28, 5, 17, 16, 1, 53]. In 3D computer vision, CroCo [45, 46] pioneered cross-view completion as a pretext task, where one view is partially masked and reconstructed using visible portions along with a second reference view of the same scene (see Fig. 1 left). This approach allows to learn powerful 3D features and has made possible impressive results on downstream tasks such as 3D reconstruction [43, 12, 47, 40], 3D pose regression [7], camera calibration [24], 4D reconstruction [50, 21, 42] and Gaussian splatting [33, 49].
Despite its success, the cross-view completion pretext task has a fundamental limitation: pretraining is only effective when training pairs have substantial overlap. For instance, Croco [45, 46] relies on pairs with at least 50% overlap. We are particularly interested in relative pose regression between two images, and using this pretext task impairs the performance of pose regression on pairs with limited overlap, as our results show.
In this paper, we present Alligat0R (see Fig. 1 right), a novel pre-training approach that reformulates the cross-view learning objective as a co-visibility segmentation task. Instead of reconstructing masked regions, our method explicitly predicts whether each pixel in one image is: (1) co-visible in the second image, (2) occluded, or (3) outside the field of view (FOV). This formulation offers several advantages: it can use image pairs with any degree of overlap, it provides interpretable predictions that reveal the model’s geometric understanding, and it aligns more directly with the correspondence reasoning required in downstream binocular vision tasks.
To enable our approach, we introduce Cub3, a large-scale dataset comprising 2.5 million image pairs with dense co-visibility annotations derived from the nuScenes dataset. This dataset includes challenging scenarios with varying degrees of overlap between views, providing a more diverse training signal than previous approaches.
One might notice that Alligat0R is not strictly self-supervised, unlike some earlier pre-training methods, as it relies on co-visibility annotations. However, the same can be said for CroCo, which requires filtering image pairs to retain only those with at least 50% overlap. The creation of our co-visibility annotations is more complex than for CroCo’s, however the process is fully automated. This process is very similar to the one used in RUBIK [26] and relies on registered video sequences and depth predictions.
Our contributions can be summarized as follows:
  1. 1.
    We introduce co-visibility segmentation as a novel pre-training objective for binocular vision tasks, replacing the cross-completion-based approach of prior work while maintaining the same architecture.
  2. 2.
    We demonstrate that co-visibility segmentation-based pretraining significantly benefits from its ability to learn from challenging image pairs, i.e. from any overlap.
  3. 3.
    We create and release Cub3, a large-scale dataset with dense co-visibility annotations derived from the nuScenes dataset.
  4. 4.
    We show that Alligat0R significantly outperforms CroCo pre-training when fine-tuned on the relative pose regression task, particularly for challenging scenarios with limited overlap between views.
  5. 5.
    We demonstrate the effectiveness of our learning based on geometric difficulty and provide insights into the model’s cross-view reasoning capabilities through interpretable segmentation outputs.
Our experiments demonstrate that explicitly learning to understand co-visibility relationships between image pairs leads to more robust and transferable features for relative pose regression compared to reconstruction-based approaches.
2 Related Work
Pretraining on Image Pairs.
To the best of our knowledge, CroCo [45] pioneered the extension of masked image modeling [18] to image pairs, introducing cross-view completion. CroCo v2 [46] applied this cross-view completion framework to large amounts of data. P-Match [54] introduced a variation where both images are partially masked to pre-train an image matching model. Another variant, masked appearance transfer, was proposed in [51] for object tracking, with a similar approach in [34]. Let us highlight that the foundational models DUSt3R [43] and MASt3R [24] are built on CroCo. While all these works rely on cross-view completion, we depart from this framework and propose a novel pretraining objective based on co-visibility segmentation.
Considering co-visibility across image pairs.
Ground-truth co-visibility masks are widely used in image matching methods [30, 10, 25, 52, 35, 44, 11, 4, 41, 37, 9, 13, 15, 23, 14, 20, 36, 31] to compute overlaps and select training pairs. Several approaches [31, 11, 38, 9, 14] also leverage these masks to learn matchability scores, which are used at test time to select correspondences. Additionally, [19] jointly learns co-visibility and relative planar pose between 360 panoramas. To the best of our knowledge, our work is the first to learn visual representations from co-visibility segmentation.
Fine-tuning for Relative Pose Regression.
Reloc3R [7] recently fine-tuned the foundational model DUSt3R [43], which is based on CroCo, on large-scale data for relative pose regression. In our experiments, we adopt a similar strategy to evaluate our novel pretraining method against CroCo. However, we focus on the autonomous driving scenario with metric relative translation, whereas Reloc3R predicts only the direction of the relative translation.
3 Method
In this section, we describe our novel pre-training approach, Alligat0R, which replaces the cross-view completion objective used in CroCo [45, 46] by a segmentation task. We first outline the overall architecture, which remains largely similar to CroCo, and then detail our co-visibility segmentation pre-training objective. Finally, we explain how our model can be fine-tuned for downstream tasks such as relative pose regression.
3.1 Architecture Overview
Figure 2: Overview of Alligat0R. (a) During pre-training, we use the same architecture as CroCo but replace the reconstruction objective with a co-visibility segmentation task, without masking, where each pixel in one view is classified as co-visible, occluded, or outside FOV with respect to the other view. (b) For fine-tuning on the relative pose regression task, we pool features from both views, process them through a shared MLP, and use separate heads for predicting translation and rotation.
Our architecture closely follows the design of CroCo, which consists of an encoder and a decoder. The encoder is a ViT [8] that processes the input images by dividing them into non-overlapping patches and encoding them into feature representations. The decoder is another transformer that combines information from both views, using cross-attention layers, to make predictions.
We adopt a symmetric architecture for our forward pass, where both images are processed in the same way without any masking. This differs from CroCo, which uses an asymmetric approach that heavily masks one image while leaving the other unmasked. Our symmetric design is not only more efficient but also better aligned with downstream binocular vision tasks, which typically process unmasked images.
As shown in Fig. 2 (a), given two images $I_{1}$ and $I_{2}$ of the same scene taken from different viewpoints, we first divide both into non-overlapping patches, denoted as tokens $p_{1}=\{p^{1}_{1},\ldots,p^{N_{1}}_{1}\}$ and $p_{2}=\{p^{1}_{2},\ldots,p^{N_{2}}_{2}\}$.
The encoder $E_{\theta}$ processes the tokens from both images independently:
$F_{1}=E_{\theta}(p_{1}),\quad F_{2}=E_{\theta}(p_{2})\>,$ (1)
and the decoder $D_{\phi}$ then takes these features and processes them to enable cross-view reasoning:
$O_{1}=D_{\phi}(F_{1},F_{2}),\quad O_{2}=D_{\phi}(F_{2},F_{1})\>,$ (2)
where $O_{1}$ and $O_{2}$ represent the output features from the decoder. The decoder contains cross-attention mechanisms that allow information exchange between features from both views.
3.2 Co-Visibility Segmentation Pre-Training
A key difference between our approach and CroCo lies in the pre-training objective. Instead of reconstructing masked pixels, we formulate the pre-training task as a co-visibility segmentation problem. Our goal is to predict for each pixel in each image whether it is:
  1. 1.
    Co-visible: the pixel corresponds to a 3D point that is also visible in the other image.
  2. 2.
    Occluded: the pixel corresponds to a 3D point that is occluded in the other image.
  3. 3.
    Outside FOV: the pixel corresponds to a 3D point that is outside the field of view in the other image.
Formally, for each patch token, the decoder produces a feature vector that is processed by a fully-connected layer to output probabilities for each of the three classes, for each pixel:
$\hat{y}_{ikj}=\text{softmax}(W\cdot O_{ik}+b)_{j}\>,$ (3)
where $\hat{y}_{ikj}\in\mathbb{R}^{3}$ represents the predicted probabilities for the three co-visibility classes for pixel $j$ in patch $k$ of image $i$, and $W$ and $b$ are learnable parameters.
During pre-training, we optimize the network using a cross-entropy loss in each image:
$\mathcal{L}_{\text{ce}_{i}}=-\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\log(\hat{y}_{ij,c_{ij}})\>,$ (4)
where $c_{ij}$ is the ground-truth class in pixel $j$ of image $i$, and $N_{i}$ is the number of pixels in each image.
This formulation offers several advantages over the cross-view-completion-based approach. First, it allows us to use image pairs with any degree of overlap, as we explicitly model cases where pixels have no co-visible region in the other view. Second, it provides interpretable outputs that directly reveal the model’s understanding of scene geometry. Third, it better aligns with downstream tasks such as pose regression, where both images are fully visible.
3.3 Fine-Tuning for Relative Pose Regression
After pre-training, we fine-tune our model for the task of relative pose regression. We add a pose regression head on top of the pre-trained encoder-decoder architecture, while keeping the original co-visibility segmentation head. This design allows the model to leverage the geometric understanding acquired during pre-training.
The pose regression head consists of several components as shown in Fig. 2 (b). First, we apply global average pooling to the decoder outputs for both images:
$f_{1}=\text{GlobalAvgPool}(O_{1}),\quad f_{2}=\text{GlobalAvgPool}(O_{2})\>.$ (5)
The pooled features are concatenated and processed through a shared MLP:
$f_{\text{shared}}=\text{MLP}([f_{1},f_{2}])\>.$ (6)
We then use separate heads for predicting the metric relative translation vector $t\in\mathbb{R}^{3}$ and the relative rotation represented as a quaternion $q\in\mathbb{R}^{4}$:
$t=\text{MLP}_{t}(f_{\text{shared}}),\quad q=\text{MLP}_{q}(f_{\text{shared}})\>.$ (7)
Our fine-tuning stage consists of two phases:
  1. 1.
    First, we freeze the pre-trained encoder, decoder and co-visibility segmentation head, and only train the pose regression head. During this phase, we use a homoscedastic loss [22] combining MSE losses for translation and quaternion predictions:
    $\mathcal{L}_{\text{pose}}=\frac{1}{2\sigma_{t}^{2}}\|t-\hat{t}\|^{2}+\frac{1}{2\sigma_{q}^{2}}\|q-\hat{q}\|^{2}+\log\sigma_{t}+\log\sigma_{q}\>,$ (8)
    where $t$ and $q$ are the ground-truth metric translation and quaternion respectively, and $\sigma_{t}$ and $\sigma_{q}$ are learnable parameters that automatically balance the two loss terms.
  2. 2.
    In the second phase, we unfreeze the backbone and the co-visibility segmentation head, and train the full network with a joint loss that combines the pose regression loss and the co-visibility segmentation loss:
    $\mathcal{L}_{\text{joint}}=\frac{1}{2\sigma_{\text{pose}}^{2}}\mathcal{L}_{\text{pose}}+\frac{1}{2\sigma_{\text{seg}}^{2}}\mathcal{L}_{\text{ce}}+\log\sigma_{\text{pose}}+\log\sigma_{\text{seg}}\>,$ (9)
    where $\sigma_{\text{pose}}$ and $\sigma_{\text{seg}}$ are learnable parameters that balance the two loss components.
This training strategy ensures that the model maintains its interpretable co-visibility segmentation capabilities while being optimized for the pose regression task.
4 Cub3: A Large-Scale Co-Visibility Dataset
To enable our co-visibility segmentation approach, we introduce Cub3, a large-scale dataset comprising 2.5 million image pairs with dense co-visibility annotations derived from the autonomous driving nuScenes dataset [2]. Cub3 will be made publicly available.
4.1 Dataset Construction
We leverage the co-visibility estimation pipeline introduced in RUBIK [26] to generate pixel-level co-visibility annotations for image pairs. In brief, this pipeline uses monocular metric depth predictions from UniDepth [29] and surface normals from Depth Anything V2 [48], combined with camera poses from COLMAP [32] reconstructions, to automatically classify each pixel as either co-visible, occluded, or outside FOV with respect to another image.
Starting from the nuScenes training set, we apply this pipeline to all possible image pairs within each scene, resulting in approximately 34 million annotated pairs. For our experiments, we created two dataset variants, each containing 2.5 million image pairs:
  • Cub3-50: Image pairs with at least 50% overlap, similar to the criterion used in CroCo [46]. This dataset provides a direct comparison point with existing cross-view completion approaches.
  • Cub3-all: Image pairs with at least 5% overlap. This dataset, contains challenging image pairs to test the robustness of methods to handle low-overlap scenarios that are common in real-world applications.
Fig. 3 shows examples or our co-visibility annotations on image pairs from Cub3, illustrating how our process classifies pixels into the three categories across varying levels of difficulty.
It is also worth noting that our annotation process inherits some limitations from the underlying monocular depth estimation models. The depth predictions may occasionally struggle with challenging scenarios such as reflective surfaces, transparent objects, or regions with complex geometry or far-away objects. Consequently, some annotations in our dataset, particularly the distinction between co-visible and occluded pixels, may contain noise. Despite these imperfections, our experiments show that the scale and diversity of Cub3 enable effective training of robust co-visibility segmentation models.
4.2 Dataset Statistics
Figure 3: Co-visibility annotation examples from Cub3. For each image pair, we show the corresponding co-visibility maps with color-coding for co-visible, occluded, and outside FOV regions. Note how our annotation process handles varying degrees of overlap and challenging viewpoint changes. Let us highlight that some annotations, particularly the distinction between co-visible and occluded pixels, may contain noise, see text for details.
Figure 4: Distributions of overlap, scale ratio, and viewpoint angle in Cub3-all and Cub3-50.
Fig. 4 illustrates the distribution of our datasets across three geometric criteria introduced in RUBIK: overlap percentage, scale ratio, and viewpoint angle. The Cub3-all dataset’s distribution is heavily skewed towards very challenging pairs, which are particularly difficult for cross-completion methods like CroCo. In contrast, our co-visibility segmentation approach is designed to benefit from such challenging cases effectively.
Given our computational resources, Cub3 is currently limited to autonomous driving scenarios from nuScenes. But such scenarios have very important applications, and given its size, variety, and challenges, Cub3 is already sufficient to validate Alligat0R—very much like the initial version of CroCo [45] was demonstrated as a ”proof-of-concept”:
  • Scale: With 2.5 million annotated pairs, Cub3 provides sufficient data to train and validate our approach.
  • Geometric variety: It covers a wide range of overlaps, scales and viewpoint angles within the driving context.
  • Challenging scenarios: It includes pairs with minimal overlap that test the limits of CroCo, as was done in the original paper.
  • Application relevance: Autonomous driving represents an important application domain with significant research interest.
Our experiments on the RUBIK benchmark, which was specifically designed to evaluate challenging autonomous driving scenarios, demonstrate the effectiveness of our approach. The promising results we obtain motivate future work to extend Cub3 to more diverse environments beyond urban driving scenes. However, even within this specific domain, our method shows significant improvements over CroCo, particularly in geometrically challenging configurations.
5 Experiments
5.1 Implementation Details
We implement Alligat0R using PyTorch and conduct all experiments on NVIDIA A100 GPUs. The model architecture follows the design described in Sec. 3.1, utilizing a ViT-based encoder and transformer decoder backbone similar to CroCo, with 24 layers for the encoder and 12 for the decoder. The input images are resized to have a maximum dimension of 512 pixels while maintaining their aspect ratio, resulting in an input resolution suitable for capturing sufficient detail while remaining computationally efficient. For pre-training, we use the AdamW optimizer with a learning rate of 1.5e-4, weight decay of 0.05, and a batch size of 32 per GPU. We employ a cosine learning rate schedule with 2 epochs of warmup and train for 25 epochs on our Cub3 datasets.
For fine-tuning on relative pose regression, we follow the two-phase approach described in Sec. 3.3. In the first phase, we freeze the backbone and train only the pose regression head for 5 epochs with a learning rate of 1e-4. In the second phase, we unfreeze the entire network and jointly train with both the pose regression loss and co-visibility segmentation loss for an additional 10 epochs with a learning rate of 5e-5.
The comprehensive architecture of Alligat0R is illustrated in Fig. 2, showing both the pre-training and fine-tuning phases. During pre-training, the model learns to segment pixels in each view as co-visible, occluded, or outside the field of view with respect to the other view. During fine-tuning, we introduce a pose regression head while maintaining the segmentation capability to leverage the geometric understanding acquired during pre-training.
5.2 Main Results
Table 1: Results on RUBIK [26] Benchmark for Relative Pose Regression. See text for details.
Pre-training Training set Backbone Finetune set $5^{\circ}$ / 2m $10^{\circ}$ / 5m AUC @ 5/10/20
CroCo Cub3-50 [Uncaptioned image] Cub3-50 19.10 28.80 15.6 / 24.7 / 33.9
CroCo Cub3-all [Uncaptioned image] Cub3-all 8.90 25.20 8.5 / 22.0 / 40.8
Alligat0R Cub3-50 [Uncaptioned image] Cub3-50 24.40 32.10 14.8 / 23.1 / 31.2
Alligat0R Cub3-all [Uncaptioned image] Cub3-all 55.30 82.30 31.1 / 54.0 / 71.8
Alligat0R Cub3-all [Uncaptioned image] Cub3-all 60.30 81.90 41.6 / 61.2 / 75.5
From scratch - [Uncaptioned image] Cub3-all 29.10 52.30 27.1 / 43.7 / 58.7
Relative Pose Regression Performance. We evaluate our proposed Alligat0R pre-training approach on the relative pose regression task and compare it with CroCo pre-training using the same architecture and fine-tuning protocol. Tab. 1 presents the results on the RUBIK [26] benchmark, which is specifically designed to evaluate performance across different geometric challenges in autonomous driving scenarios.
As shown in Tab. 1, Alligat0R consistently outperforms CroCo across different training and evaluation configurations. Notably, when pre-trained on Cub3-all and fine-tuned with the backbone frozen, Alligat0R achieves significantly higher performance and reaches state-of-the-art on the RUBIK benchmark (55.30% success rate at $5^{\circ}$/2m and 82.30% at $10^{\circ}$/5m) compared to the CroCo model pre-trained on the same dataset (which only achieves 8.90% and 25.20% respectively in the frozen backbone configuration), confirming the results in the original CroCo paper. This substantial improvement demonstrates the effectiveness of our co-visibility segmentation pre-training approach, especially for challenging scenarios with varying degrees of overlap.
Impact of Training Data Distribution. Our results reveal an interesting pattern regarding the impact of training data distribution. When trained on pairs with at least 50% overlap, CroCo achieves modest performance (19.10% accuracy at $5^{\circ}$/2m and 28.80% at $10^{\circ}$/5m). However, when training on the full distribution, Alligat0R shows remarkable improvement (from 24.40% to 55.30% at $5^{\circ}$/2m and from 32.10% to 82.30% at $10^{\circ}$/5m). This confirms our hypothesis that explicitly modeling co-visibility through segmentation enables more effective learning from challenging image pairs with limited overlap.
Figure 5: Performance of CroCo and Alligat0R across different geometric challenges. Results show accuracy at the $5^{\circ}$/2m threshold for models trained on different datasets (Cub3-50 or Cub3-all) and for frozen backbones. Alligat0R trained on Cub3-all consistently outperforms other configurations, particularly for challenging cases with low overlap, large scale differences, and extreme viewpoint changes.
Fig. 5 further illustrates this advantage by breaking down performance across different overlap percentages, scale ratios and viewpoint angles. While both methods perform well on high-overlap pairs (80-100%), Alligat0R trained on Cub3-all maintains strong performance even as overlap decreases, whereas CroCo’s accuracy drops dramatically for pairs with less than 40% overlap. This demonstrates Alligat0R’s ability to generalize across a wide range of geometric configurations, which is crucial for real-world applications where overlap cannot be guaranteed.
Fine-tuning Strategy Comparison. We also compare different fine-tuning strategies, specifically the impact of freezing versus unfreezing the backbone during fine-tuning. For Alligat0R, unfreezing the backbone during fine-tuning leads to even better performance and yielding the best results (60.30% at $5^{\circ}$/2m and 81.90% at $10^{\circ}$/5m compared to 55.30% and 82.30% with the frozen backbone), suggesting that our pre-training approach learns transferable features that can be effectively adapted to the pose regression task.
Comparison with Training from Scratch. To isolate the impact of our pre-training approach, we compare the performance of a model trained from scratch for relative pose regression versus those initialized with Alligat0R or CroCo pre-training. The final row in Tab. 1 shows that training from scratch achieves notably lower performance (29.10% at $5^{\circ}$/2m and 52.30% at $10^{\circ}$/5m) compared to our pre-trained models trained on the same dataset (Cub3-all), highlighting the value of the representations learned during pre-training.
Table 2: RUBIK Benchmark. The results are reported from [26]. Success rate (in %) for each method across individual geometric criterion bins. Best and second-best values for each column are shown in bold and underlined respectively.
Overlap (%) Scale Ratio Viewpoint Angle (°) Whole Time
80–100 60–80 40–60 20–40 5–20 1.0–1.5 1.5–2.5 2.5–4.0 4.0–6.0 0–30 30–60 60–120 120–180 Dataset (ms)
Detector-based methods
ALIKED+LightGlue [52] 53.4 95.8 68.2 38.0 12.7 62.0 31.0 13.1 1.6 50.6 46.0 28.3 2.0 36.8 45
DISK+LightGlue [39] 54.2 91.4 65.9 38.7 11.8 60.4 30.8 11.6 2.4 50.3 43.8 27.4 2.7 35.9 69
SP+LightGlue [6] 64.8 93.3 68.0 36.4 10.9 61.2 28.4 12.5 1.4 49.9 43.0 28.2 0.9 35.7 43
SIFT+LightGlue [27] 68.2 92.1 61.4 32.3 9.9 57.3 26.9 9.6 1.7 49.8 39.7 23.7 0.5 33.1 194
DeDoDe v2 [10] 89.8 93.3 54.8 26.7 7.9 60.4 16.3 3.2 0.9 49.3 35.4 19.9 0.3 30.4 282
XFeat [30] 85.4 67.4 24.3 5.2 0.9 32.1 2.4 0.1 0.0 34.4 8.3 7.1 0.0 14.2 54
XFeat* [30] 62.4 69.1 27.6 7.6 1.5 32.8 4.5 0.6 0.0 33.8 9.4 9.2 0.0 15.1 82
XFeat+LighterGlue [30] 64.6 91.7 59.1 26.2 8.1 56.6 20.9 4.6 0.2 48.0 33.4 21.4 1.2 30.1 43
Detector-free methods
LoFTR [35] 87.2 88.4 47.2 17.5 5.0 51.6 10.1 2.3 0.6 43.2 27.9 15.1 0.0 24.9 185
ELoFTR [44] 56.4 90.3 50.8 22.1 6.3 51.2 15.6 4.4 0.7 42.2 30.8 18.2 0.1 26.6 124
ASpanFormer [4] 72.2 72.3 44.5 21.9 7.4 46.0 14.9 6.9 1.6 42.5 27.2 16.0 0.1 24.8 108
RoMa [11] 67.0 98.3 84.5 52.7 20.2 71.2 43.2 26.6 8.3 57.5 56.2 44.1 3.0 47.3 614
DUSt3R [43] 81.8 97.4 90.8 58.4 30.4 73.3 57.9 40.1 9.9 67.4 55.3 50.0 35.2 54.8 257
MASt3R [24] 52.0 97.5 89.6 61.0 28.4 71.2 52.3 42.5 13.8 53.5 65.6 54.5 14.1 53.6 173
Relative pose regression
Alligat0R (Ours) 100.0 89.7 80.2 61.5 44.3 78.4 61.3 43.9 23.2 77.5 62.7 51.3 37.3 60.3 57
5.3 Qualitative Analysis
Co-visibility and Reconstruction Visualization. Fig. 6 shows qualitative examples of the co-visibility segmentation and CroCo reconstructions produced by our models on pairs from RUBIK. Alligat0R successfully identifies co-visible regions, occluded areas, and regions outside FOV across varying degrees of overlap and viewpoint changes. These visualizations provide insights into the geometric understanding acquired by our model. Conversely, the best CroCo model succeeds in reconstructing masked regions for pairs with high overlaps, but completely fails on more difficult pairs.
Figure 6: Qualitative comparison between CroCo trained on Cub3-50 and Alligat0R. For each row from left to right: reference image, target image, masked target image, CroCo reconstruction, and Alligat0R segmentations (co-visible, occluded, outside FOV). CroCo struggles with reconstructing regions with limited overlap, while Alligat0R often correctly identifies co-visible, occluded, and outside FOV regions across varying degrees of overlap and viewpoint changes.
5.4 Comparison with state-of-the-art on RUBIK
In Tab. 2, we evaluate the performance of Alligat0R fine-tuned for metric relative pose estimation against state-of-the-art methods on RUBIK [26]. Let us highlight that:
  • Alligat0R was trained on nuScenes (on scenes that are not part of RUBIK),
  • whereas none of the state-of-the-art methods used nuScenes as training set.
This gives an advantage to Alligat0R over the other methods. However, RUBIK allows the evaluation over large ranges of three geometric criteria—overlap, scale ratio and viewpoint angle—which we found interesting. Despite this advantage, we argue that Alligat0R’s performance remains remarkable, as pretraining with CroCo, in contrast, leads to a significant performance drop (see Tab. 1).
With an overall success rate of 60.3%, Alligat0R ranks first. Notably, it significantly outperforms all other methods on the hardest bin across the three geometric criteria, which confirms the advantage of its pretraining on difficult image pairs. While Alligat0R’s training set was highly skewed toward challenging pairs (see Fig. 4), one might expect a performance drop on easier bins. However, this does not happen: for instance, Alligat0R still ranks first for small scale ratios and small viewpoint angles. Finally, Alligat0R is among the fastest methods, as it directly regresses the relative pose rather than producing intermediate correspondences.
5.5 Future Directions
Our results demonstrate the effectiveness of co-visibility segmentation for pre-training. From this idea, we believe that several promising directions can be explored:
  1. 1.
    Segmentation Formulations: Investigating different ways to formulate the segmentation task, including binary vs. three-class approaches.
  2. 2.
    Curriculum Learning: Experimenting with curriculums that gradually increase the difficulty of training examples to see if this could further improve performance.
  3. 3.
    Multi-task Learning: Studying whether jointly training on multiple downstream tasks can leverage the geometric understanding provided by our co-visibility segmentation approach.
  4. 4.
    Diverse Environments: Extending Alligat0R to more diverse indoor and outdoor environments beyond autonomous driving scenarios.
6 Conclusion and Future Work
In this paper, we introduced Alligat0R, a novel pre-training approach for binocular vision that reformulates cross-view completion as a co-visibility segmentation task. Unlike CroCo that relies on masked reconstruction, our approach explicitly models co-visibility relationships between image pairs, enabling more effective learning from challenging examples with limited overlap.
We demonstrated the effectiveness of our approach on the relative pose regression task in the autonomous driving scenario, where Alligat0R significantly outperforms CroCo, particularly on challenging scenes with limited overlap between views. This improvement is especially notable using a diverse training set that include pairs with varying degrees of overlap, highlighting the ability of our approach to learn and benefit from geometrically challenging examples.
To enable our approach, we created Cub3, a large-scale dataset of 2.5 million image pairs with dense co-visibility annotations derived from the nuScenes dataset. This dataset provides a valuable resource for future research in binocular vision tasks and will be made publicly available.
Our work demonstrates that explicitly modeling co-visibility through segmentation provides a more effective and interpretable approach to binocular vision tasks compared to reconstruction-based methods, paving the way for more robust visual perception systems.
Supplementary Material
7 Supplementary Material
In this supplementary material, we provide additional details about our Alligat0R model, including training curves and visualizations that didn’t fit in the main paper but may be of interest to readers.
7.1 Pre-training Learning Curves
Figure 7 shows the learning curves for the pre-training phase of both CroCo and Alligat0R on the Cub3-50 and Cub3-all datasets. As can be observed, Alligat0R’s loss converges smoothly on both datasets, indicating stable training. There are fewer number of steps for CroCo as we have a batch size of 64 per GPU, compared to Alligat0R where we have a batch size of 32.
Figure 7: Learning curves during pre-training for CroCo (reconstruction loss) and Alligat0R (segmentation loss) on Cub3-50 and Cub3-all datasets. Alligat0R shows stable convergence on both datasets, while CroCo exhibits instability on the more challenging Cub3-all dataset which contains image pairs with lower overlap.
7.2 Fine-tuning Learning Curves
Figure 8 presents the fine-tuning curves for the pose regression task. The plots show the pose loss during training. Alligat0R fine-tuned on Cub3-all shows faster convergence and reaches higher accuracy than the same configuration for CroCo, highlighting the transferability of features learned through co-visibility segmentation on difficult pairs.
Figure 8: Learning curves during fine-tuning for pose regression. Alligat0R pre-trained on Cub3-all converges faster and achieves higher success rates than other methods, demonstrating the effectiveness of our co-visibility segmentation pre-training approach.
7.3 Detailed Performance
Figure 9 provides a more detailed breakdown of performance across different geometric criteria than what was presented in the main paper. This visualization further emphasizes Alligat0R’s strong performance across all difficulty ranges, especially on very difficult scenarios where traditional methods struggle.
Figure 9: Detailed breakdown of performance across different geometric criteria. Success rate either for R@5° or t@2m (bottom-left and top-right of each triangle, respectively), when projecting results onto individual geometric criteria of RUBIK. For each method, we show three plots corresponding to the projection over overlap (left), scale ratio (middle), and viewpoint angle (right)

### v2

@@
-Relative Camera Pose Regression
+Thibaut Loiseau1    Guillaume Bourmaud2    Vincent Lepetit1
 Abstract
 Pre-training techniques have greatly advanced computer vision, with CroCo’s cross-view completion approach yielding impressive results in tasks like 3D reconstruction and pose regression.
-However, this method requires substantial overlap between training pairs, limiting its effectiveness.
-We introduce Alligat0R, a novel pre-training approach that reformulates cross-view learning as a co-visibility segmentation task.
-Our method predicts whether each pixel in one image is co-visible in the second image, occluded, or outside the field of view (FOV), enabling the use of image pairs with any degree of overlap and providing interpretable predictions.
-To support this, we present Cub3, a large-scale dataset with 2.5 million image pairs and dense co-visibility annotations derived from the nuScenes dataset.
-This dataset includes diverse scenarios with varying degrees of overlap.
-The experiments show that Alligat0R significantly outperforms CroCo in relative pose regression, especially in scenarios with limited overlap.
-Alligat0R and Cub3 will be made publicly available.
-Figure 1: We introduce Alligat0R, a pretraining method for binocular vision robust to wide baselines.
-Alligat0R explicitly segments pixels as co-visible, occluded, or outside field-of-view, overcoming the fundamental limitation of CroCo [45, 46] which attempts to reconstruct potentially unobservable regions.
-We demonstrate Alligat0R on relative pose regression and show that it allows us to significantly outperform the state-of-the-art, especially on pairs with very low overlap.
+However, cross-view completion is ill-posed in non-covisible regions, limiting its effectiveness.
+We introduce Alligat0R, a novel pre-training approach that replaces cross-view learning with a covisibility segmentation task.
+Our method predicts whether each pixel in one image is covisible in the second image, occluded, or outside the field of view, making the pre-training effective in both covisible and non-covisible regions, and provides interpretable predictions.
+To support this, we present Cub3, a large-scale dataset with 5M image pairs and dense covisibility annotations derived from the nuScenes and ScanNet datasets.
+Cub3 includes diverse scenarios with varying degrees of overlap.
+The experiments show that our novel pre-training method Alligat0R significantly outperforms CroCo in relative pose regression.
+Code is available at https://github.com/thibautloiseau/alligat0r.
+Figure 1: We introduce Alligat0R, a novel pretraining method for binocular vision.
+Alligat0R explicitly segments pixels as covisible, occluded, or outside field-of-view, overcoming the fundamental limitation of CroCo [weinzaepfel2022croco, weinzaepfel2023croco] which attempts to reconstruct potentially non-covisible regions.
 1 Introduction
-Pre-training techniques have revolutionized computer vision by enabling large models to learn rich representations [3, 28, 5, 17, 16, 1, 53].
-In 3D computer vision, CroCo [45, 46] pioneered cross-view completion as a pretext task, where one view is partially masked and reconstructed using visible portions along with a second reference view of the same scene (see Fig.
-1 left).
-This approach allows to learn powerful 3D features and has made possible impressive results on downstream tasks such as 3D reconstruction [43, 12, 47, 40], 3D pose regression [7], camera calibration [24], 4D reconstruction [50, 21, 42] and Gaussian splatting [33, 49].
-Despite its success, the cross-view completion pretext task has a fundamental limitation: pretraining is only effective when training pairs have substantial overlap.
-For instance, Croco [45, 46] relies on pairs with at least 50% overlap.
-We are particularly interested in relative pose regression between two images, and using this pretext task impairs the performance of pose regression on pairs with limited overlap, as our results show.
-In this paper, we present Alligat0R (see Fig.
-1 right), a novel pre-training approach that reformulates the cross-view learning objective as a co-visibility segmentation task.
-Instead of reconstructing masked regions, our method explicitly predicts whether each pixel in one image is: (1) co-visible in the second image, (2) occluded, or (3) outside the field of view (FOV).
-This formulation offers several advantages: it can use image pairs with any degree of overlap, it provides interpretable predictions that reveal the model’s geometric understanding, and it aligns more directly with the correspondence reasoning required in downstream binocular vision tasks.
-To enable our approach, we introduce Cub3, a large-scale dataset comprising 2.5 million image pairs with dense co-visibility annotations derived from the nuScenes dataset.
-This dataset includes challenging scenarios with varying degrees of overlap between views, providing a more diverse training signal than previous approaches.
-One might notice that Alligat0R is not strictly self-supervised, unlike some earlier pre-training methods, as it relies on co-visibility annotations.
+Pre-training techniques have revolutionized computer vision by enabling large models to learn rich representations [caron2021emerging, oquab2024dinov2, chen2020simple, he2022masked, grill2020bootstrap, bao2022beit, zhou2022image].
+In 3D computer vision, CroCo [weinzaepfel2022croco, weinzaepfel2023croco] pioneered cross-view completion as a pretext task, where one view is partially masked and reconstructed using visible portions along with a second reference view of the same scene (see Figure˜1 left).
+This approach allows to learn powerful 3D features and has made possible impressive results on downstream tasks such as 3D reconstruction [wang2024dust3r, elflein2025light3r, yang2025fast3r, wang20243d], 3D pose regression [dong2024reloc3r], camera calibration [leroy2024grounding], 4D reconstruction [zhang2024monst3r, jin2024stereo4d, wang2025continuous] and Gaussian splatting [smart2024splatt3r, ye2024no].
+Despite its success, the cross-view completion pretext task has a fundamental limitation: pretraining is only effective in covisible regions, as the task is ill-posed in non-covisible regions (see Figure˜1).
+As a consequence, Croco [weinzaepfel2022croco, weinzaepfel2023croco] relies on pairs with at least 50% of covisible regions.
+In this paper, we present Alligat0R, a novel pre-training approach that offers an alternative to the cross-view completion objective through a covisibility segmentation task.
+Instead of reconstructing masked regions, our method explicitly predicts whether each pixel in one image is: (1) covisible in the second image, (2) occluded, or (3) outside the field of view (FOV).
+This formulation offers several advantages: it makes the pre-training effective in both covisible and non-covisible regions, it provides interpretable predictions that reveal the model’s geometric understanding, and it aligns more directly with the correspondence reasoning required in downstream binocular vision tasks.
+To enable our approach, we introduce Cub3, a large-scale dataset comprising two sub datasets, each of 2.5 million image pairs with dense covisibility annotations derived from nuScenes [caesar2020nuscenes] and ScanNet [dai2017scannet] datasets.
+Cub3 includes challenging scenarios with varying degrees of overlap between views, providing a more diverse training signal than previous approaches.
+One might notice that Alligat0R is not strictly self-supervised, unlike some earlier pre-training methods, as it relies on covisibility annotations.
 However, the same can be said for CroCo, which requires filtering image pairs to retain only those with at least 50% overlap.
-The creation of our co-visibility annotations is more complex than for CroCo’s, however the process is fully automated.
-This process is very similar to the one used in RUBIK [26] and relies on registered video sequences and depth predictions.
+The creation of our covisibility annotations for nuScenes is more complex than for ScanNet, as the ground truth poses and depth maps are not available, however the process is fully automated.
+This process is very similar to the one used in RUBIK [loiseau2025rubik] and relies on registered video sequences and depth predictions.
 Our contributions can be summarized as follows:
   1. 1.
-    We introduce co-visibility segmentation as a novel pre-training objective for binocular vision tasks, replacing the cross-completion-based approach of prior work while maintaining the same architecture.
+    We introduce covisibility segmentation as a novel pre-training objective for binocular vision tasks, replacing the cross-completion-based approach of prior work while maintaining the same network architecture.
   2. 2.
-    We demonstrate that co-visibility segmentation-based pretraining significantly benefits from its ability to learn from challenging image pairs, i.e. from any overlap.
+    We create and release Cub3, a large-scale dataset with dense covisibility annotations derived from nuScenes and ScanNet datasets.
   3. 3.
-    We create and release Cub3, a large-scale dataset with dense co-visibility annotations derived from the nuScenes dataset.
+    We show that Alligat0R provides an effective alternative to CroCo pre-training when fine-tuned on the relative pose regression task, particularly for challenging scenarios with limited overlap between views.
   4. 4.
-    We show that Alligat0R significantly outperforms CroCo pre-training when fine-tuned on the relative pose regression task, particularly for challenging scenarios with limited overlap between views.
-  5. 5.
-    We demonstrate the effectiveness of our learning based on geometric difficulty and provide insights into the model’s cross-view reasoning capabilities through interpretable segmentation outputs.
-Our experiments demonstrate that explicitly learning to understand co-visibility relationships between image pairs leads to more robust and transferable features for relative pose regression compared to reconstruction-based approaches.
+    We demonstrate the effectiveness of our pre-training and provide insights into the model’s cross-view reasoning capabilities through interpretable segmentation outputs.
+Our experiments demonstrate that explicitly learning to understand covisibility relationships between image pairs leads to more robust and transferable features for relative pose regression compared to reconstruction-based approaches.
 2 Related Work
 Pretraining on Image Pairs.
-To the best of our knowledge, CroCo [45] pioneered the extension of masked image modeling [18] to image pairs, introducing cross-view completion.
-CroCo v2 [46] applied this cross-view completion framework to large amounts of data.
-P-Match [54] introduced a variation where both images are partially masked to pre-train an image matching model.
-Another variant, masked appearance transfer, was proposed in [51] for object tracking, with a similar approach in [34].
-Let us highlight that the foundational models DUSt3R [43] and MASt3R [24] are built on CroCo.
-While all these works rely on cross-view completion, we depart from this framework and propose a novel pretraining objective based on co-visibility segmentation.
-Considering co-visibility across image pairs.
-Ground-truth co-visibility masks are widely used in image matching methods [30, 10, 25, 52, 35, 44, 11, 4, 41, 37, 9, 13, 15, 23, 14, 20, 36, 31] to compute overlaps and select training pairs.
-Several approaches [31, 11, 38, 9, 14] also leverage these masks to learn matchability scores, which are used at test time to select correspondences.
-Additionally, [19] jointly learns co-visibility and relative planar pose between 360 panoramas.
-To the best of our knowledge, our work is the first to learn visual representations from co-visibility segmentation.
+To the best of our knowledge, CroCo [weinzaepfel2022croco] pioneered the extension of masked image modeling [hondru2024masked] to image pairs, introducing cross-view completion.
+CroCo v2 [weinzaepfel2023croco] applied this cross-view completion framework to large amounts of data.
+P-Match [zhu2023pmatch] introduced a variation where both images are partially masked to pre-train an image matching model.
+Another variant, masked appearance transfer, was proposed in [zhao2023representation] for object tracking, with a similar approach in [song2023compact].
+Let us highlight that the foundational models DUSt3R [wang2024dust3r] and MASt3R [leroy2024grounding] are built on CroCo.
+While all these works rely on cross-view completion, we depart from this framework and propose a novel pretraining objective based on covisibility segmentation.
+Considering covisibility across image pairs.
+Ground-truth covisibility masks are widely used in image matching methods [potje2024xfeat, edstedt2024dedode, lindenberger2023lightglue, zhao2023aliked, sun2021loftr, wang2024efficient, edstedt2024roma, chen2022aspanformer, wang2022matchformer, tangquadtree, edstedt2023dkm, fan2023occ, gleize2023silk, kim2025learning, germainvisual, jiang2021cotr, tan2022eco, sarlin2020superglue] to compute overlaps and select training pairs.
+Several approaches [sarlin2020superglue, edstedt2024roma, truong2023pdc, edstedt2023dkm, germainvisual, bono2023end] also leverage these masks to learn matchability scores, which are used at test time to select correspondences.
+Additionally, [hutchcroft2022covispose] jointly learns covisibility and relative planar pose between 360 panoramas.
+To the best of our knowledge, our work is the first to learn visual representations from covisibility segmentation.
 Fine-tuning for Relative Pose Regression.
-Reloc3R [7] recently fine-tuned the foundational model DUSt3R [43], which is based on CroCo, on large-scale data for relative pose regression.
+Reloc3R [dong2024reloc3r] recently fine-tuned the foundational model DUSt3R [wang2024dust3r], which is based on CroCo, on large-scale data for relative pose regression.
 In our experiments, we adopt a similar strategy to evaluate our novel pretraining method against CroCo.
-However, we focus on the autonomous driving scenario with metric relative translation, whereas Reloc3R predicts only the direction of the relative translation.
+However, we fine-tune Alligat0R for metric relative translation, whereas Reloc3R predicts only the direction of the relative translation.
 3 Method
-In this section, we describe our novel pre-training approach, Alligat0R, which replaces the cross-view completion objective used in CroCo [45, 46] by a segmentation task.
-We first outline the overall architecture, which remains largely similar to CroCo, and then detail our co-visibility segmentation pre-training objective.
+In this section, we describe our novel pre-training approach, Alligat0R, which replaces the cross-view completion objective used in CroCo [weinzaepfel2022croco, weinzaepfel2023croco] by a segmentation task.
+We first outline the overall architecture, which remains largely similar to CroCo, and then detail our covisibility segmentation pre-training objective.
 Finally, we explain how our model can be fine-tuned for downstream tasks such as relative pose regression.
@@
 Figure 2: Overview of Alligat0R.
-(a) During pre-training, we use the same architecture as CroCo but replace the reconstruction objective with a co-visibility segmentation task, without masking, where each pixel in one view is classified as co-visible, occluded, or outside FOV with respect to the other view.
+(a) During pre-training, we use the same architecture as CroCo but replace the reconstruction objective with a covisibility segmentation task, without masking, where each pixel in one view is classified as covisible, occluded, or outside FOV with respect to the other view.
 (b) For fine-tuning on the relative pose regression task, we pool features from both views, process them through a shared MLP, and use separate heads for predicting translation and rotation.
 Our architecture closely follows the design of CroCo, which consists of an encoder and a decoder.
-The encoder is a ViT [8] that processes the input images by dividing them into non-overlapping patches and encoding them into feature representations.
+The encoder is a ViT [dosovitskiy2020image] that processes the input images by dividing them into non-overlapping patches and encoding them into feature representations.
 The decoder is another transformer that combines information from both views, using cross-attention layers, to make predictions.
@@
 Our symmetric design is not only more efficient but also better aligned with downstream binocular vision tasks, which typically process unmasked images.
-As shown in Fig.
-2 (a), given two images $I_{1}$ and $I_{2}$ of the same scene taken from different viewpoints, we first divide both into non-overlapping patches, denoted as tokens $p_{1}=\{p^{1}_{1},\ldots,p^{N_{1}}_{1}\}$ and $p_{2}=\{p^{1}_{2},\ldots,p^{N_{2}}_{2}\}$.
+As shown in Figure˜2 (a), given two images $\mathtt{I}_{1}$ and $\mathtt{I}_{2}$ of the same scene taken from different viewpoints, we first divide both into non-overlapping patches, denoted as tokens $\mathtt{P}_{1}=\{\mathtt{P}^{1}_{1},\ldots,\mathtt{P}^{N_{1}}_{1}\}$ and $\mathtt{P}_{2}=\{\mathtt{P}^{1}_{2},\ldots,\mathtt{P}^{N_{2}}_{2}\}$.
 The encoder $E_{\theta}$ processes the tokens from both images independently:
-$F_{1}=E_{\theta}(p_{1}),\quad F_{2}=E_{\theta}(p_{2})\>,$ (1)
+$\mathtt{F}_{1}=E_{\theta}(\mathtt{P}_{1}),\quad\mathtt{F}_{2}=E_{\theta}(\mathtt{P}_{2})\>,$ (1)
 and the decoder $D_{\phi}$ then takes these features and processes them to enable cross-view reasoning:
-$O_{1}=D_{\phi}(F_{1},F_{2}),\quad O_{2}=D_{\phi}(F_{2},F_{1})\>,$ (2)
-where $O_{1}$ and $O_{2}$ represent the output features from the decoder. The decoder contains cross-attention mechanisms that allow information exchange between features from both views.
-3.2 Co-Visibility Segmentation Pre-Training
+$\mathtt{O}_{1}=D_{\phi}(\mathtt{F}_{1},\mathtt{F}_{2}),\quad\mathtt{O}_{2}=D_{\phi}(\mathtt{F}_{2},\mathtt{F}_{1})\>,$ (2)
+where $\mathtt{O}_{1}$ and $\mathtt{O}_{2}$ represent the output features from the decoder.
+The decoder contains cross-attention mechanisms that allow information exchange between features from both views.
+3.2 Covisibility Segmentation Pre-Training
 A key difference between our approach and CroCo lies in the pre-training objective.
-Instead of reconstructing masked pixels, we formulate the pre-training task as a co-visibility segmentation problem.
+Instead of reconstructing masked pixels, we formulate the pre-training task as a covisibility segmentation problem.
 Our goal is to predict for each pixel in each image whether it is:
   1. 1.
-    Co-visible: the pixel corresponds to a 3D point that is also visible in the other image.
+    Covisible: the pixel corresponds to a 3D point that is also visible in the other image.
   2. 2.
@@
 Formally, for each patch token, the decoder produces a feature vector that is processed by a fully-connected layer to output probabilities for each of the three classes, for each pixel:
-$\hat{y}_{ikj}=\text{softmax}(W\cdot O_{ik}+b)_{j}\>,$ (3)
-where $\hat{y}_{ikj}\in\mathbb{R}^{3}$ represents the predicted probabilities for the three co-visibility classes for pixel $j$ in patch $k$ of image $i$, and $W$ and $b$ are learnable parameters.
+$\hat{\mathtt{Y}}_{ikj}=\text{softmax}(\mathtt{W}\cdot\mathtt{O}_{ik}+\mathbf{b})_{j}\>,$ (3)
+where $\hat{\mathtt{Y}}_{ikj}\in\mathbb{R}^{3}$ represents the predicted probabilities for the three covisibility classes for pixel $j$ in patch $k$ of image $i$, and $\mathtt{W}$ and $\mathbf{b}$ are learnable parameters.
 During pre-training, we optimize the network using a cross-entropy loss in each image:
-$\mathcal{L}_{\text{ce}_{i}}=-\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\log(\hat{y}_{ij,c_{ij}})\>,$ (4)
-where $c_{ij}$ is the ground-truth class in pixel $j$ of image $i$, and $N_{i}$ is the number of pixels in each image.
+$\mathcal{L}_{\text{ce}_{i}}=-\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\ln(\hat{\mathtt{Y}}_{ij,c_{ij}})\>,$ (4)
+where $\hat{\mathtt{Y}}_{ij,c_{ij}}$ is the predicted probability for the ground-truth class $c_{ij}$ in pixel $j$ of image $i$, and $N_{i}$ is the number of pixels in each image.
+The total cross-entropy loss is the sum over both images: $\mathcal{L}_{\text{ce}}=\mathcal{L}_{\text{ce}_{1}}+\mathcal{L}_{\text{ce}_{2}}$.
 This formulation offers several advantages over the cross-view-completion-based approach.
-First, it allows us to use image pairs with any degree of overlap, as we explicitly model cases where pixels have no co-visible region in the other view.
+First, it makes the pre-training effective in both covisible and non-covisible regions, as we explicitly model cases where pixels have no covisible region in the other view.
 Second, it provides interpretable outputs that directly reveal the model’s understanding of scene geometry.
@@
 After pre-training, we fine-tune our model for the task of relative pose regression.
-We add a pose regression head on top of the pre-trained encoder-decoder architecture, while keeping the original co-visibility segmentation head.
+We add a pose regression head on top of the pre-trained encoder-decoder architecture, while keeping the original covisibility segmentation head.
 This design allows the model to leverage the geometric understanding acquired during pre-training.
-The pose regression head consists of several components as shown in Fig. 2 (b). First, we apply global average pooling to the decoder outputs for both images:
-$f_{1}=\text{GlobalAvgPool}(O_{1}),\quad f_{2}=\text{GlobalAvgPool}(O_{2})\>.$ (5)
+The pose regression head consists of several components as shown in Figure˜2 (b). First, we apply global average pooling to the decoder outputs for both images:
+$\mathbf{f}_{1}=\text{GlobalAvgPool}(\mathtt{O}_{1}),\quad\mathbf{f}_{2}=\text{GlobalAvgPool}(\mathtt{O}_{2})\>.$ (5)
 The pooled features are concatenated and processed through a shared MLP:
-$f_{\text{shared}}=\text{MLP}([f_{1},f_{2}])\>.$ (6)
-We then use separate heads for predicting the metric relative translation vector $t\in\mathbb{R}^{3}$ and the relative rotation represented as a quaternion $q\in\mathbb{R}^{4}$:
-$t=\text{MLP}_{t}(f_{\text{shared}}),\quad q=\text{MLP}_{q}(f_{\text{shared}})\>.$ (7)
+$\mathbf{f}_{\text{shared}}=\text{MLP}([\mathbf{f}_{1},\mathbf{f}_{2}])\>.$ (6)
+We then use separate heads for predicting the metric relative translation vector $\hat{\mathbf{t}}\in\mathbb{R}^{3}$ and the relative rotation represented as a quaternion $\hat{\mathbf{q}}\in\mathbb{R}^{4}$:
+$\hat{\mathbf{t}}=\text{MLP}_{\mathbf{t}}(\mathbf{f}_{\text{shared}}),\quad\hat{\mathbf{q}}=\text{MLP}_{\mathbf{q}}(\mathbf{f}_{\text{shared}})\>.$ (7)
+We use unit quaternions with L2 normalization, treating them as 4D vectors in the homoscedastic loss function.
 Our fine-tuning stage consists of two phases:
   1. 1.
-    First, we freeze the pre-trained encoder, decoder and co-visibility segmentation head, and only train the pose regression head.
-During this phase, we use a homoscedastic loss [22] combining MSE losses for translation and quaternion predictions:
-    $\mathcal{L}_{\text{pose}}=\frac{1}{2\sigma_{t}^{2}}\|t-\hat{t}\|^{2}+\frac{1}{2\sigma_{q}^{2}}\|q-\hat{q}\|^{2}+\log\sigma_{t}+\log\sigma_{q}\>,$ (8)
-    where $t$ and $q$ are the ground-truth metric translation and quaternion respectively, and $\sigma_{t}$ and $\sigma_{q}$ are learnable parameters that automatically balance the two loss terms.
+    First, we freeze the pre-trained encoder, decoder and covisibility segmentation head, and only train the pose regression head.
+During this phase, we use a homoscedastic loss [kendall2018multi] combining MSE losses for translation and quaternion predictions:
+    $\mathcal{L}_{\text{pose}}=\frac{1}{2\sigma_{\mathbf{t}}^{2}}\|\mathbf{t}-\hat{\mathbf{t}}\|^{2}+\frac{1}{2\sigma_{\mathbf{q}}^{2}}\|\mathbf{q}-\hat{\mathbf{q}}\|^{2}+\log\sigma_{\mathbf{t}}+\log\sigma_{\mathbf{q}}\>,$ (8)
+    where $\mathbf{t}$ and $\mathbf{q}$ are the ground-truth metric translation and quaternion respectively, and $\sigma_{\mathbf{t}}$ and $\sigma_{\mathbf{q}}$ are learnable parameters that automatically balance the two loss terms.
   2. 2.
-    In the second phase, we unfreeze the backbone and the co-visibility segmentation head, and train the full network with a joint loss that combines the pose regression loss and the co-visibility segmentation loss:
-    $\mathcal{L}_{\text{joint}}=\frac{1}{2\sigma_{\text{pose}}^{2}}\mathcal{L}_{\text{pose}}+\frac{1}{2\sigma_{\text{seg}}^{2}}\mathcal{L}_{\text{ce}}+\log\sigma_{\text{pose}}+\log\sigma_{\text{seg}}\>,$ (9)
-    where $\sigma_{\text{pose}}$ and $\sigma_{\text{seg}}$ are learnable parameters that balance the two loss components.
-This training strategy ensures that the model maintains its interpretable co-visibility segmentation capabilities while being optimized for the pose regression task.
-4 Cub3: A Large-Scale Co-Visibility Dataset
-To enable our co-visibility segmentation approach, we introduce Cub3, a large-scale dataset comprising 2.5 million image pairs with dense co-visibility annotations derived from the autonomous driving nuScenes dataset [2].
+    In the second phase, we unfreeze the backbone and the covisibility segmentation head, and train the full network with a joint loss that combines the pose regression loss and the covisibility segmentation loss:
+    $\mathcal{L}_{\text{joint}}=\mathcal{L}_{\text{pose}}+\frac{1}{2\sigma_{\text{seg}}^{2}}\mathcal{L}_{\text{ce}}+\log\sigma_{\text{seg}}\>,$ (9)
+    where $\sigma_{\text{seg}}$ is a learnable parameter.
+This training strategy ensures that the model maintains its interpretable covisibility segmentation capabilities while being optimized for the pose regression task.
+4 Cub3: A Large-Scale Covisibility Dataset
+To enable our covisibility segmentation approach, we introduce Cub3, a large-scale dataset comprising two sub datasets each of 2.5 million image pairs with dense covisibility annotations derived from both the autonomous driving nuScenes dataset [caesar2020nuscenes], and the indoor ScanNet [dai2017scannet] dataset.
 Cub3 will be made publicly available.
 4.1 Dataset Construction
-We leverage the co-visibility estimation pipeline introduced in RUBIK [26] to generate pixel-level co-visibility annotations for image pairs.
-In brief, this pipeline uses monocular metric depth predictions from UniDepth [29] and surface normals from Depth Anything V2 [48], combined with camera poses from COLMAP [32] reconstructions, to automatically classify each pixel as either co-visible, occluded, or outside FOV with respect to another image.
-Starting from the nuScenes training set, we apply this pipeline to all possible image pairs within each scene, resulting in approximately 34 million annotated pairs.
-For our experiments, we created two dataset variants, each containing 2.5 million image pairs:
-  • Cub3-50: Image pairs with at least 50% overlap, similar to the criterion used in CroCo [46]. This dataset provides a direct comparison point with existing cross-view completion approaches.
+nuScenes – We leverage the covisibility estimation pipeline introduced in RUBIK [loiseau2025rubik] to generate pixel-level covisibility annotations for image pairs from nuScenes.
+In brief, this pipeline uses monocular metric depth predictions from UniDepth [piccinelli2024unidepth] and surface normals from Depth Anything V2 [yang2024depth], combined with camera poses from COLMAP [schonberger2016structure] reconstructions, to automatically classify each pixel as either covisible, occluded, or outside FOV with respect to another image.
+We apply this pipeline to all possible image pairs within each scene, resulting in approximately 34 million annotated pairs.
+ScanNet – The pipeline is similar but simpler for ScanNet, as we rely on the ground truth depths and camera poses to get all covisibility annotations.
+For our experiments, we created two dataset variants, each containing 5M image pairs (2.5M from nuScenes and 2.5M from ScanNet):
+  • Cub3-50: Image pairs with at least 50% overlap, similar to the criterion used in CroCo [weinzaepfel2023croco].
+This dataset provides a direct comparison point with existing cross-view completion approaches.
   • Cub3-all: Image pairs with at least 5% overlap.
 This dataset, contains challenging image pairs to test the robustness of methods to handle low-overlap scenarios that are common in real-world applications.
-Fig. 3 shows examples or our co-visibility annotations on image pairs from Cub3, illustrating how our process classifies pixels into the three categories across varying levels of difficulty.
-It is also worth noting that our annotation process inherits some limitations from the underlying monocular depth estimation models.
+Figure˜3 shows examples or our covisibility annotations on image pairs from Cub3, illustrating how our process classifies pixels into the three categories across varying levels of difficulty.
+It is also worth noting that our annotation process, especially on nuScenes, inherits some limitations from the underlying monocular depth estimation models.
 The depth predictions may occasionally struggle with challenging scenarios such as reflective surfaces, transparent objects, or regions with complex geometry or far-away objects.
-Consequently, some annotations in our dataset, particularly the distinction between co-visible and occluded pixels, may contain noise.
-Despite these imperfections, our experiments show that the scale and diversity of Cub3 enable effective training of robust co-visibility segmentation models.
+Consequently, some annotations in our dataset, particularly the distinction between covisible and occluded pixels, may contain noise.
+Despite these imperfections, our experiments show that the scale and diversity of Cub3 enable effective training of robust covisibility segmentation models.
 4.2 Dataset Statistics
-Figure 3: Co-visibility annotation examples from Cub3.
-For each image pair, we show the corresponding co-visibility maps with color-coding for co-visible, occluded, and outside FOV regions.
+Figure 3: Covisibility annotation examples from Cub3 for nuScenes (top) and ScanNet (bottom).
+For each image pair, we show the corresponding covisibility maps with color-coding for covisible, occluded, and outside FOV regions.
 Note how our annotation process handles varying degrees of overlap and challenging viewpoint changes.
-Let us highlight that some annotations, particularly the distinction between co-visible and occluded pixels, may contain noise, see text for details.
-Figure 4: Distributions of overlap, scale ratio, and viewpoint angle in Cub3-all and Cub3-50.
-Fig.
-4 illustrates the distribution of our datasets across three geometric criteria introduced in RUBIK: overlap percentage, scale ratio, and viewpoint angle.
+Let us highlight that some annotations, particularly the distinction between covisible and occluded pixels, may contain noise, especially for nuScenes, and we demonstrate in the experiments that Alligat0R is highly robust to this noise.
+Figure 4: Distributions of overlap, scale ratio, and viewpoint angle in Cub3-all and Cub3-50 for nuScenes (left) and ScanNet (right).
+Figure˜4 illustrates the distribution of our datasets across three geometric criteria introduced in RUBIK [loiseau2025rubik]: overlap percentage, scale ratio, and viewpoint angle.
 The Cub3-all dataset’s distribution is heavily skewed towards very challenging pairs, which are particularly difficult for cross-completion methods like CroCo.
-In contrast, our co-visibility segmentation approach is designed to benefit from such challenging cases effectively.
-Given our computational resources, Cub3 is currently limited to autonomous driving scenarios from nuScenes.
-But such scenarios have very important applications, and given its size, variety, and challenges, Cub3 is already sufficient to validate Alligat0R—very much like the initial version of CroCo [45] was demonstrated as a ”proof-of-concept”:
-  • Scale: With 2.5 million annotated pairs, Cub3 provides sufficient data to train and validate our approach.
-  • Geometric variety: It covers a wide range of overlaps, scales and viewpoint angles within the driving context.
+In contrast, our covisibility segmentation approach is designed to benefit from such challenging cases effectively.
+Given our computational resources, Cub3 is currently limited to either autonomous driving scenarios from nuScenes, or indoor scenes from ScanNet.
+But such scenarios have very important applications, and given its size, variety, and challenges, Cub3 is already sufficient to validate Alligat0R—very much like the initial version of CroCo [weinzaepfel2022croco] was demonstrated as a "proof-of-concept":
+  • Scale: With 5M annotated pairs, Cub3 provides sufficient data to train and validate our approach on both outdoor and indoor scenarios.
+  • Geometric variety: It covers a wide range of overlaps, scales and viewpoint angles within the driving and indoor contexts.
   • Challenging scenarios: It includes pairs with minimal overlap that test the limits of CroCo, as was done in the original paper.
-  • Application relevance: Autonomous driving represents an important application domain with significant research interest.
-Our experiments on the RUBIK benchmark, which was specifically designed to evaluate challenging autonomous driving scenarios, demonstrate the effectiveness of our approach.
-The promising results we obtain motivate future work to extend Cub3 to more diverse environments beyond urban driving scenes.
+Our experiments on the RUBIK benchmark, which was specifically designed to evaluate challenging autonomous driving scenarios, as well as results on ScanNet1500 test set demonstrate the effectiveness of our approach.
+The promising results we obtain motivate future work to extend Cub3 to more diverse environments beyond urban driving scenes and indoor scenes.
 However, even within this specific domain, our method shows significant improvements over CroCo, particularly in geometrically challenging configurations.
@@
 We implement Alligat0R using PyTorch and conduct all experiments on NVIDIA A100 GPUs.
-The model architecture follows the design described in Sec.
-3.1, utilizing a ViT-based encoder and transformer decoder backbone similar to CroCo, with 24 layers for the encoder and 12 for the decoder.
-The input images are resized to have a maximum dimension of 512 pixels while maintaining their aspect ratio, resulting in an input resolution suitable for capturing sufficient detail while remaining computationally efficient.
-For pre-training, we use the AdamW optimizer with a learning rate of 1.5e-4, weight decay of 0.05, and a batch size of 32 per GPU.
-We employ a cosine learning rate schedule with 2 epochs of warmup and train for 25 epochs on our Cub3 datasets.
-For fine-tuning on relative pose regression, we follow the two-phase approach described in Sec.
-3.3.
-In the first phase, we freeze the backbone and train only the pose regression head for 5 epochs with a learning rate of 1e-4.
-In the second phase, we unfreeze the entire network and jointly train with both the pose regression loss and co-visibility segmentation loss for an additional 10 epochs with a learning rate of 5e-5.
-The comprehensive architecture of Alligat0R is illustrated in Fig.
-2, showing both the pre-training and fine-tuning phases.
-During pre-training, the model learns to segment pixels in each view as co-visible, occluded, or outside the field of view with respect to the other view.
-During fine-tuning, we introduce a pose regression head while maintaining the segmentation capability to leverage the geometric understanding acquired during pre-training.
+The model architecture follows the design described in Section˜3.1.
+The input images are classically resized to have a maximum dimension of 512 pixels while maintaining their aspect ratio.
+More information is provided in the supplementary material.
 5.2 Main Results
-Table 1: Results on RUBIK [26] Benchmark for Relative Pose Regression. See text for details.
-Pre-training Training set Backbone Finetune set $5^{\circ}$ / 2m $10^{\circ}$ / 5m AUC @ 5/10/20
-CroCo Cub3-50 [Uncaptioned image] Cub3-50 19.10 28.80 15.6 / 24.7 / 33.9
-CroCo Cub3-all [Uncaptioned image] Cub3-all 8.90 25.20 8.5 / 22.0 / 40.8
-Alligat0R Cub3-50 [Uncaptioned image] Cub3-50 24.40 32.10 14.8 / 23.1 / 31.2
-Alligat0R Cub3-all [Uncaptioned image] Cub3-all 55.30 82.30 31.1 / 54.0 / 71.8
-Alligat0R Cub3-all [Uncaptioned image] Cub3-all 60.30 81.90 41.6 / 61.2 / 75.5
-From scratch - [Uncaptioned image] Cub3-all 29.10 52.30 27.1 / 43.7 / 58.7
-Relative Pose Regression Performance.
-We evaluate our proposed Alligat0R pre-training approach on the relative pose regression task and compare it with CroCo pre-training using the same architecture and fine-tuning protocol.
-Tab.
-1 presents the results on the RUBIK [26] benchmark, which is specifically designed to evaluate performance across different geometric challenges in autonomous driving scenarios.
-As shown in Tab.
-1, Alligat0R consistently outperforms CroCo across different training and evaluation configurations.
-Notably, when pre-trained on Cub3-all and fine-tuned with the backbone frozen, Alligat0R achieves significantly higher performance and reaches state-of-the-art on the RUBIK benchmark (55.30% success rate at $5^{\circ}$/2m and 82.30% at $10^{\circ}$/5m) compared to the CroCo model pre-trained on the same dataset (which only achieves 8.90% and 25.20% respectively in the frozen backbone configuration), confirming the results in the original CroCo paper.
-This substantial improvement demonstrates the effectiveness of our co-visibility segmentation pre-training approach, especially for challenging scenarios with varying degrees of overlap.
+Table 1: Results on RUBIK [loiseau2025rubik] and ScanNet1500 [dai2017scannet] for metric relative pose regression. For all experiments, fine-tuning is performed using Cub3-all.
+RUBIK ScanNet1500 Pre-training Pre-Training Set Backbone $5^{\circ}$ / 0.5m $5^{\circ}$ / 2m $10^{\circ}$ / 5m $10^{\circ}$ / 0.25m $10^{\circ}$ / 0.5m $10^{\circ}$ / 1m CroCo Cub3-50 [Uncaptioned image] 4.4 21.8 48.1 48.3 64.1 69.3 CroCo Cub3-all [Uncaptioned image] 2.4 8.9 25.2 8.7 18.3 24.9 Alligat0R Cub3-50 [Uncaptioned image] 9.3 32.4 58.4 47.5 59.9 63.4 Alligat0R Cub3-all [Uncaptioned image] 24.0 55.3 82.3 78.2 90.1 92.5 CroCo Cub3-50 [Uncaptioned image] 12.4 38.3 66.7 75.7 87.4 91.5 CroCo Cub3-all [Uncaptioned image] 12.5 38.6 66.2 64.1 79.1 85.0 Alligat0R Cub3-50 [Uncaptioned image] 21.2 53.3 78.1 82.8 91.7 93.9 Alligat0R Cub3-all [Uncaptioned image] 24.6 60.3 81.9 85.5 92.5 95.1 From scratch Cub3-all [Uncaptioned image] 10.8 29.1 52.3 37.0 51.8 57.2
+Metric Relative Pose Regression Performance.
+We evaluate our proposed Alligat0R pre-training approach on the metric relative pose regression task and compare it with CroCo pre-training using the same architecture and fine-tuning protocol.
+Table˜1 presents the results on the RUBIK [loiseau2025rubik] benchmark, which is specifically designed to evaluate performance across different geometric challenges in autonomous driving scenarios, and on the indoor ScanNet1500 [dai2017scannet] benchmark.
+As shown in Table˜1, Alligat0R consistently outperforms CroCo across different training and evaluation configurations.
+Notably, when pre-trained on Cub3-all and fine-tuned with the backbone frozen, Alligat0R achieves significantly higher performance and reaches state-of-the-art on the RUBIK benchmark (55.3% success rate at $5^{\circ}$/2m and 82.3% at $10^{\circ}$/5m) compared to the CroCo model pre-trained on the same dataset (which only achieves 8.9% and 25.2% respectively in the frozen backbone configuration).
+This substantial improvement demonstrates the effectiveness of our covisibility segmentation pre-training approach, especially for challenging scenarios with varying degrees of overlap.
 Impact of Training Data Distribution.
-Our results reveal an interesting pattern regarding the impact of training data distribution.
-When trained on pairs with at least 50% overlap, CroCo achieves modest performance (19.10% accuracy at $5^{\circ}$/2m and 28.80% at $10^{\circ}$/5m).
-However, when training on the full distribution, Alligat0R shows remarkable improvement (from 24.40% to 55.30% at $5^{\circ}$/2m and from 32.10% to 82.30% at $10^{\circ}$/5m).
-This confirms our hypothesis that explicitly modeling co-visibility through segmentation enables more effective learning from challenging image pairs with limited overlap.
-Figure 5: Performance of CroCo and Alligat0R across different geometric challenges.
+A key observation is the contrasting behavior of Alligat0R and CroCo when trained on different data distributions.
+While CroCo performs similarly or better when trained on easier pairs (Cub3-50) compared to harder pairs (Cub3-all), Alligat0R shows a different trend.
+Our method systematically improves when trained on the full distribution (Cub3-all).
+This demonstrates that Alligat0R can effectively learn from challenging examples, while CroCo struggles to benefit from them.
+Figure 5: Performance of CroCo and Alligat0R across different geometric challenges on RUBIK.
 Results show accuracy at the $5^{\circ}$/2m threshold for models trained on different datasets (Cub3-50 or Cub3-all) and for frozen backbones.
 Alligat0R trained on Cub3-all consistently outperforms other configurations, particularly for challenging cases with low overlap, large scale differences, and extreme viewpoint changes.
-Fig.
-5 further illustrates this advantage by breaking down performance across different overlap percentages, scale ratios and viewpoint angles.
+Figure˜5 further illustrates this advantage by breaking down performance across different overlap percentages, scale ratios and viewpoint angles.
 While both methods perform well on high-overlap pairs (80-100%), Alligat0R trained on Cub3-all maintains strong performance even as overlap decreases, whereas CroCo’s accuracy drops dramatically for pairs with less than 40% overlap.
@@
 We also compare different fine-tuning strategies, specifically the impact of freezing versus unfreezing the backbone during fine-tuning.
-For Alligat0R, unfreezing the backbone during fine-tuning leads to even better performance and yielding the best results (60.30% at $5^{\circ}$/2m and 81.90% at $10^{\circ}$/5m compared to 55.30% and 82.30% with the frozen backbone), suggesting that our pre-training approach learns transferable features that can be effectively adapted to the pose regression task.
+One can see that Alligat0R trained on Cub3-all is already quite high with a frozen backbone and always outperforms CroCo even when its backbone is unfrozen.
+This suggests that our pre-training strategy is better aligned with the downstream pose regression task.
 Comparison with Training from Scratch.
 To isolate the impact of our pre-training approach, we compare the performance of a model trained from scratch for relative pose regression versus those initialized with Alligat0R or CroCo pre-training.
-The final row in Tab.
-1 shows that training from scratch achieves notably lower performance (29.10% at $5^{\circ}$/2m and 52.30% at $10^{\circ}$/5m) compared to our pre-trained models trained on the same dataset (Cub3-all), highlighting the value of the representations learned during pre-training.
+The bottom row of Table˜1 shows that training from scratch achieves notably lower performance, highlighting the value of the representations learned during pre-training.
+Robustness to Label Noise.
+To address concerns about sensitivity to covisibility segmentation errors, we conducted a sensitivity analysis by training Alligat0R on ScanNet with 20% random label noise injected during pre-training.
+Surprisingly, this actually improved performance across all thresholds (+2.2%, +2.4%, +0.8% respectively), as shown in Table˜2.
+This counterintuitive result suggests that label noise acts as a form of regularization similar to label smoothing [szegedy2016rethinking], encouraging learning of more generalizable features.
+Importantly, this demonstrates that our approach remains robust even when covisibility annotations are imperfect.
+Table 2: Sensitivity analysis: Impact of label noise on Alligat0R performance on ScanNet1500.
+Results show that 20% random label noise during pre-training actually improves performance, demonstrating robustness to annotation errors.
+ScanNet1500
+Method Noise $10^{\circ}$ / 0.25m $10^{\circ}$ / 0.5m $10^{\circ}$ / 1m
+Alligat0R 0% 85.5 92.5 95.1
+Alligat0R 20% 87.7 94.9 95.9
+CroCo N/A 75.7 87.4 91.5
 Table 3: RUBIK Benchmark.
-The results are reported from [26].
+The results are reported from [loiseau2025rubik].
 Success rate (in %) for each method across individual geometric criterion bins.
@@
 80–100 60–80 40–60 20–40 5–20 1.0–1.5 1.5–2.5 2.5–4.0 4.0–6.0 0–30 30–60 60–120 120–180 Dataset (ms)
-Detector-based methods
-ALIKED+LightGlue [52] 53.4 95.8 68.2 38.0 12.7 62.0 31.0 13.1 1.6 50.6 46.0 28.3 2.0 36.8 45
-DISK+LightGlue [39] 54.2 91.4 65.9 38.7 11.8 60.4 30.8 11.6 2.4 50.3 43.8 27.4 2.7 35.9 69
-SP+LightGlue [6] 64.8 93.3 68.0 36.4 10.9 61.2 28.4 12.5 1.4 49.9 43.0 28.2 0.9 35.7 43
-SIFT+LightGlue [27] 68.2 92.1 61.4 32.3 9.9 57.3 26.9 9.6 1.7 49.8 39.7 23.7 0.5 33.1 194
-DeDoDe v2 [10] 89.8 93.3 54.8 26.7 7.9 60.4 16.3 3.2 0.9 49.3 35.4 19.9 0.3 30.4 282
-XFeat [30] 85.4 67.4 24.3 5.2 0.9 32.1 2.4 0.1 0.0 34.4 8.3 7.1 0.0 14.2 54
-XFeat* [30] 62.4 69.1 27.6 7.6 1.5 32.8 4.5 0.6 0.0 33.8 9.4 9.2 0.0 15.1 82
-XFeat+LighterGlue [30] 64.6 91.7 59.1 26.2 8.1 56.6 20.9 4.6 0.2 48.0 33.4 21.4 1.2 30.1 43
 Detector-free methods
-LoFTR [35] 87.2 88.4 47.2 17.5 5.0 51.6 10.1 2.3 0.6 43.2 27.9 15.1 0.0 24.9 185
-ELoFTR [44] 56.4 90.3 50.8 22.1 6.3 51.2 15.6 4.4 0.7 42.2 30.8 18.2 0.1 26.6 124
-ASpanFormer [4] 72.2 72.3 44.5 21.9 7.4 46.0 14.9 6.9 1.6 42.5 27.2 16.0 0.1 24.8 108
-RoMa [11] 67.0 98.3 84.5 52.7 20.2 71.2 43.2 26.6 8.3 57.5 56.2 44.1 3.0 47.3 614
-DUSt3R [43] 81.8 97.4 90.8 58.4 30.4 73.3 57.9 40.1 9.9 67.4 55.3 50.0 35.2 54.8 257
-MASt3R [24] 52.0 97.5 89.6 61.0 28.4 71.2 52.3 42.5 13.8 53.5 65.6 54.5 14.1 53.6 173
+LoFTR [sun2021loftr] 87.2 88.4 47.2 17.5 5.0 51.6 10.1 2.3 0.6 43.2 27.9 15.1 0.0 24.9 185
+ELoFTR [wang2024efficient] 56.4 90.3 50.8 22.1 6.3 51.2 15.6 4.4 0.7 42.2 30.8 18.2 0.1 26.6 124
+ASpanFormer [chen2022aspanformer] 72.2 72.3 44.5 21.9 7.4 46.0 14.9 6.9 1.6 42.5 27.2 16.0 0.1 24.8 108
+RoMa [edstedt2024roma] 67.0 98.3 84.5 52.7 20.2 71.2 43.2 26.6 8.3 57.5 56.2 44.1 3.0 47.3 614
+DUSt3R [wang2024dust3r] 81.8 97.4 90.8 58.4 30.4 73.3 57.9 40.1 9.9 67.4 55.3 50.0 35.2 54.8 257
+MASt3R [leroy2024grounding] 52.0 97.5 89.6 61.0 28.4 71.2 52.3 42.5 13.8 53.5 65.6 54.5 14.1 53.6 173
 Relative pose regression
@@
 5.3 Qualitative Analysis
-Co-visibility and Reconstruction Visualization.
-Fig.
-6 shows qualitative examples of the co-visibility segmentation and CroCo reconstructions produced by our models on pairs from RUBIK.
-Alligat0R successfully identifies co-visible regions, occluded areas, and regions outside FOV across varying degrees of overlap and viewpoint changes.
+Covisibility and Reconstruction Visualization.
+Figure˜6 shows qualitative examples of the covisibility segmentation and CroCo reconstructions produced by our models on pairs from Cub3.
+Alligat0R successfully identifies covisible regions, occluded areas, and regions outside FOV across varying degrees of overlap and viewpoint changes.
 These visualizations provide insights into the geometric understanding acquired by our model.
-Conversely, the best CroCo model succeeds in reconstructing masked regions for pairs with high overlaps, but completely fails on more difficult pairs.
-Figure 6: Qualitative comparison between CroCo trained on Cub3-50 and Alligat0R.
-For each row from left to right: reference image, target image, masked target image, CroCo reconstruction, and Alligat0R segmentations (co-visible, occluded, outside FOV).
-CroCo struggles with reconstructing regions with limited overlap, while Alligat0R often correctly identifies co-visible, occluded, and outside FOV regions across varying degrees of overlap and viewpoint changes.
+Conversely, while Croco succeeds in reconstructing masked covisible regions, the reconstructions are blurred in masked non-covisible regions as cross-completion is ill-posed in such areas.
+ Reference    Target       Masked   CroCo recons.   Alligat0R segmentations
+Figure 6: Qualitative comparison.
+CroCo’s reconstructions are blurred in masked non-covisible regions, while Alligat0R often correctly identifies covisible, occluded, outside FOV regions across varying degrees of overlap and viewpoint changes.
 5.4 Comparison with state-of-the-art on RUBIK
-In Tab. 2, we evaluate the performance of Alligat0R fine-tuned for metric relative pose estimation against state-of-the-art methods on RUBIK [26]. Let us highlight that:
+In Table˜3, we evaluate the performance of Alligat0R fine-tuned for metric relative pose estimation against state-of-the-art methods on RUBIK [loiseau2025rubik]. Let us highlight that:
   • Alligat0R was trained on nuScenes (on scenes that are not part of RUBIK),
@@
 However, RUBIK allows the evaluation over large ranges of three geometric criteria—overlap, scale ratio and viewpoint angle—which we found interesting.
-Despite this advantage, we argue that Alligat0R’s performance remains remarkable, as pretraining with CroCo, in contrast, leads to a significant performance drop (see Tab.
-1).
+Despite this advantage, we argue that Alligat0R’s performance remains remarkable, as pretraining with CroCo, in contrast, leads to a significant performance drop (see Table˜1).
 With an overall success rate of 60.3%, Alligat0R ranks first.
 Notably, it significantly outperforms all other methods on the hardest bin across the three geometric criteria, which confirms the advantage of its pretraining on difficult image pairs.
-While Alligat0R’s training set was highly skewed toward challenging pairs (see Fig.
-4), one might expect a performance drop on easier bins.
+While Alligat0R’s training set was highly skewed toward challenging pairs (see Figure˜4), one might expect a performance drop on easier bins.
 However, this does not happen: for instance, Alligat0R still ranks first for small scale ratios and small viewpoint angles.
 Finally, Alligat0R is among the fastest methods, as it directly regresses the relative pose rather than producing intermediate correspondences.
-5.5 Future Directions
-Our results demonstrate the effectiveness of co-visibility segmentation for pre-training. From this idea, we believe that several promising directions can be explored:
-  1. 1.
-    Segmentation Formulations: Investigating different ways to formulate the segmentation task, including binary vs. three-class approaches.
-  2. 2.
-    Curriculum Learning: Experimenting with curriculums that gradually increase the difficulty of training examples to see if this could further improve performance.
-  3. 3.
-    Multi-task Learning: Studying whether jointly training on multiple downstream tasks can leverage the geometric understanding provided by our co-visibility segmentation approach.
-  4. 4.
-    Diverse Environments: Extending Alligat0R to more diverse indoor and outdoor environments beyond autonomous driving scenarios.
+5.5 Out-of-Domain Generalization
+To demonstrate the generalization capabilities of our method beyond the training domains, we evaluate Alligat0R on zero-shot correspondence estimation using the ETH3D dataset [schops2017multi].
+We use the correlations from decoder features to estimate correspondences and measure performance using Average EndPoint Error (AEPE) as described in [an2025cross].
+Zero-shot Correspondence Estimation.
+As detailed in the supplementary material, Alligat0R consistently outperforms CroCo when trained on the same datasets, and even surpasses CroCo v2 which was pre-trained on 5 datasets.
+This demonstrates that our covisibility segmentation pre-training learns more generalizable features for dense correspondence tasks.
+Comparison with Official CroCo v2.
+We also compare against the official CroCo v2 model, fine-tuned on our Cub3-all dataset for pose regression.
+As shown in Table˜4, Alligat0R significantly outperforms CroCo v2 across all metrics, demonstrating the effectiveness of our covisibility segmentation approach even when compared to a model pre-trained on multiple datasets.
+Table 4: Comparison with official CroCo v2 fine-tuned on Cub3-all dataset for pose regression on RUBIK benchmark. Alligat0R significantly outperforms CroCo v2 across all metrics.
+RUBIK
+Method $5^{\circ}$ / 0.5m $5^{\circ}$ / 2m $10^{\circ}$ / 5m
+Alligat0R (Frozen) 24.0 55.3 82.3
+CroCo v2 (Frozen) 5.8 15.8 34.1
+Alligat0R (Unfrozen) 24.6 60.3 81.9
+CroCo v2 (Unfrozen) 14.6 44.5 73.1
+6 Limitations
+While our approach demonstrates significant improvements over CroCo, it has several limitations:
+  • First, our current evaluation is limited to urban driving scenes from nuScenes and indoor scenes from ScanNet.
+The performance of Alligat0R on more diverse environments, such as natural scenes or extreme weather conditions, remains to be investigated.
+  • Second, our pre-training method requires ground truth depth maps and poses to produce dense covisibility annotations.
+While we investigated a fully automated pseudo ground truth generation for nuScenes, such a pre-processing is computationally intensive.
+This dependency on 3D data might limit the scalability of our approach to scenarios where such data is not readily available.
+  • Finally, while we show promising results on metric relative pose regression, the effectiveness of our pre-training approach for other binocular vision tasks, such as stereo matching or optical flow estimation, remains to be explored.
 7 Conclusion and Future Work
-In this paper, we introduced Alligat0R, a novel pre-training approach for binocular vision that reformulates cross-view completion as a co-visibility segmentation task.
-Unlike CroCo that relies on masked reconstruction, our approach explicitly models co-visibility relationships between image pairs, enabling more effective learning from challenging examples with limited overlap.
-We demonstrated the effectiveness of our approach on the relative pose regression task in the autonomous driving scenario, where Alligat0R significantly outperforms CroCo, particularly on challenging scenes with limited overlap between views.
-This improvement is especially notable using a diverse training set that include pairs with varying degrees of overlap, highlighting the ability of our approach to learn and benefit from geometrically challenging examples.
-To enable our approach, we created Cub3, a large-scale dataset of 2.5 million image pairs with dense co-visibility annotations derived from the nuScenes dataset.
-This dataset provides a valuable resource for future research in binocular vision tasks and will be made publicly available.
-Our work demonstrates that explicitly modeling co-visibility through segmentation provides a more effective and interpretable approach to binocular vision tasks compared to reconstruction-based methods, paving the way for more robust visual perception systems.
-Supplementary Material
-7 Supplementary Material
-In this supplementary material, we provide additional details about our Alligat0R model, including training curves and visualizations that didn’t fit in the main paper but may be of interest to readers.
-7.1 Pre-training Learning Curves
+We introduced Alligat0R, a novel pre-training approach that replaces cross-view completion with covisibility segmentation.
+Unlike CroCo’s masked reconstruction, our approach explicitly models covisibility relationships, enabling more effective learning from challenging geometric configurations.
+We demonstrated significant improvements over CroCo on metric relative pose regression and zero-shot correspondence estimation, particularly on challenging scenes with limited overlap.
+Our approach’s robustness to geometrically challenging examples and superior generalization capabilities highlight the effectiveness of covisibility segmentation for binocular vision tasks.
+We created Cub3, a large-scale dataset of 5M image pairs with dense covisibility annotations from nuScenes and ScanNet.
+Our work demonstrates that explicitly modeling covisibility provides an effective alternative to binocular vision tasks compared to reconstruction-based methods.
+Acknowledgments
+This project has received funding from the Bosch Research Foundation (Bosch Forschungsstiftung), the European Union (ERC Advanced Grant explorer Funding ID #101097259), and was granted access to the HPC resources of IDRIS under the allocation 2024-AD011014905R1 made by GENCI.
+Appendix A Supplementary Material
+This supplementary material provides additional details about our Alligat0R model, including training curves, implementation details, and visualizations that complement the main paper.
+A.1 Pre-training Learning Curves
 Figure 7 shows the learning curves for the pre-training phase of both CroCo and Alligat0R on the Cub3-50 and Cub3-all datasets.
-As can be observed, Alligat0R’s loss converges smoothly on both datasets, indicating stable training.
-There are fewer number of steps for CroCo as we have a batch size of 64 per GPU, compared to Alligat0R where we have a batch size of 32.
-Figure 7: Learning curves during pre-training for CroCo (reconstruction loss) and Alligat0R (segmentation loss) on Cub3-50 and Cub3-all datasets.
-Alligat0R shows stable convergence on both datasets, while CroCo exhibits instability on the more challenging Cub3-all dataset which contains image pairs with lower overlap.
-7.2 Fine-tuning Learning Curves
+Alligat0R’s loss converges smoothly on both datasets, indicating stable training.
+Figure 7: Learning curves during pre-training for CroCo (reconstruction loss) and Alligat0R (segmentation loss) on Cub3-50 and Cub3-all datasets for nuScenes (left) and ScanNet (right).
+Alligat0R shows stable convergence on both datasets.
+A.2 Fine-tuning Learning Curves
 Figure 8 presents the fine-tuning curves for the pose regression task.
-The plots show the pose loss during training.
-Alligat0R fine-tuned on Cub3-all shows faster convergence and reaches higher accuracy than the same configuration for CroCo, highlighting the transferability of features learned through co-visibility segmentation on difficult pairs.
-Figure 8: Learning curves during fine-tuning for pose regression.
-Alligat0R pre-trained on Cub3-all converges faster and achieves higher success rates than other methods, demonstrating the effectiveness of our co-visibility segmentation pre-training approach.
-7.3 Detailed Performance
-Figure 9 provides a more detailed breakdown of performance across different geometric criteria than what was presented in the main paper.
-This visualization further emphasizes Alligat0R’s strong performance across all difficulty ranges, especially on very difficult scenarios where traditional methods struggle.
+The plots show the full loss during training (Eq.
+9 in the main paper for Alligat0R, Eq.
+8 for CroCo).
+Alligat0R fine-tuned on Cub3-all shows faster convergence and reaches higher accuracy than the same configuration for CroCo, highlighting the transferability of features learned through covisibility segmentation on difficult pairs.
+The loss increase at 5 epochs corresponds to unfreezing the backbone.
+For Alligat0R, at 5 epochs we also add the segmentation loss to maintain interpretability.
+Figure 8: Learning curves during fine-tuning for pose regression for nuScenes (left) and ScanNet (right).
+Alligat0R pre-trained on Cub3-all converges faster and achieves higher success rates than other methods, demonstrating the effectiveness of our covisibility segmentation pre-training approach.
+A.3 Detailed Performance Analysis
+Figure 9 provides a comprehensive breakdown of performance across different geometric criteria on RUBIK.
+This visualization emphasizes Alligat0R’s strong performance across all difficulty ranges, particularly in challenging scenarios where traditional methods struggle.
 Figure 9: Detailed breakdown of performance across different geometric criteria.
 Success rate either for R@5° or t@2m (bottom-left and top-right of each triangle, respectively), when projecting results onto individual geometric criteria of RUBIK.
-For each method, we show three plots corresponding to the projection over overlap (left), scale ratio (middle), and viewpoint angle (right)
+For each method, we show three plots corresponding to the projection over overlap (left), scale ratio (middle), and viewpoint angle (right).
+A.4 Implementation Details
+We use a ViT-based encoder and transformer decoder backbone similar to CroCo, with 24 layers for the encoder and 12 for the decoder.
+For pre-training, we use the AdamW optimizer with a learning rate of 1.5e-4, weight decay of 0.05, and a batch size of 32 per GPU.
+We employ a cosine learning rate schedule with 2 epochs of warmup and train for 25 epochs on our Cub3 datasets.
+For fine-tuning on relative pose regression, we follow the two-phase approach described in Section 3.3 of the main paper:
+  • Phase 1: Freeze the backbone and train only the pose regression head for 5 epochs with a learning rate of 1e-4
+  • Phase 2: Unfreeze the entire network and jointly train with both the pose regression loss and covisibility segmentation loss for an additional 10 epochs with a learning rate of 5e-5
+The comprehensive architecture of Alligat0R is illustrated in Figure 2 of the main paper, showing both the pre-training and fine-tuning phases.
+During pre-training, the model learns to segment pixels in each view as covisible, occluded, or outside the field of view with respect to the other view.
+During fine-tuning, we introduce a pose regression head while maintaining the segmentation capability to leverage the geometric understanding acquired during pre-training.
+A.5 Data Splits and Sampling Protocol
+To ensure proper evaluation and avoid data leakage, we carefully separate training and test data:
+Data Splits:
+  • RUBIK benchmark: Uses test scenes from nuScenes, while our Cub3 dataset uses only the training split of nuScenes, with completely disjoint scenes.
+  • ScanNet1500 benchmark: Derived from the standard ScanNet test split, whereas Cub3 uses scenes from the ScanNet training split exclusively.
+Sampling Protocol: We pre-filter all possible pairs from all training scenes with at least 5% overlap for the "all" version of Cub3, and at least 50% for the "50" version.
+We then sample from all those pre-filtered pairs to extract the desired number of samples and covisibility masks.
+The RUBIK benchmark creation follows the protocol described in [loiseau2025rubik] and is extracted from test scenes, while ScanNet1500 was created from test scenes of ScanNet as first introduced in [sarlin2020superglue].
+A.6 Additional Ablation Studies
+A.6.1 Impact of the Number of Classes
+We implemented two variants of Alligat0R to investigate the importance of our pre-training with the three classes (covisible, occluded, outside FOV).
+We tried pre-training with only two classes, either covisible or not (in this case occluded and outside FOV are merged), or inside FOV or not (in this case, covisible and occluded are merged) on Cub3-all for nuScenes.
+The results on the RUBIK benchmark are presented in Table˜5.
+Table 5: Results on RUBIK for metric relative pose regression when pre-training with only two classes. For all experiments, pre-training and fine-tuning is performed using Cub3-all.
+RUBIK
+Classes $5^{\circ}$ / 0.5m $5^{\circ}$ / 2m $10^{\circ}$ / 5m
+Covisible or not 22.5 55.7 80.0
+Inside FOV or not 24.6 59.6 81.9
+All 3 classes 24.6 60.3 81.9
+While the performance improvement with three classes is modest, we believe that the model’s knowledge about occluded regions could be beneficial for other tasks.
+As noted in the main paper, the annotated maps may contain noise between covisible and occluded zones due to reliance on monocular depth predictions.
+A.6.2 Non-metric Relative Pose Regression
+We investigated whether our pre-training is useful for non-metric relative pose regression (regressing only the angle for translation) and compared our results with CroCo pre-training on the ScanNet1500 benchmark.
+For this experiment, we changed our pose regression head using the one from Reloc3r, along with its loss function.
+The results are shown in Table˜6.
+Table 6: Results on ScanNet1500 for non-metric relative pose regression.
+For Alligat0R, pre-training and fine-tuning is performed using Cub3-all, whereas for CroCo, pre-training is performed using Cub3-50 and fine-tuning using Cub3-all.
+ScanNet1500
+Pre-Training AUC@5 AUC@10 AUC@20
+CroCo 13.2 34.1 57.1
+Alligat0R 20.5 43.9 66.2
+Alligat0R significantly outperforms CroCo on this task, demonstrating the versatility of our pre-training approach.
+A.7 Zero-shot Correspondence Estimation on ETH3D
+To demonstrate the generalization capabilities of our method beyond the training domains, we evaluate Alligat0R on zero-shot correspondence estimation using the ETH3D dataset [schops2017multi].
+We use the correlations from decoder features to estimate correspondences and measure performance using Average EndPoint Error (AEPE) as described in [an2025cross].
+Table 7: Zero-shot correspondence estimation on ETH3D dataset using AEPE. Alligat0R consistently outperforms CroCo variants, demonstrating better generalization to out-of-domain dense matching tasks.
+Method Training Data AEPE ($\downarrow$)
+Alligat0R nuScenes-all 43.82
+CroCo nuScenes-all 77.61
+Alligat0R nuScenes-50 44.12
+CroCo nuScenes-50 92.98
+Alligat0R ScanNet-all 36.07
+CroCo ScanNet-all 56.70
+Alligat0R ScanNet-50 38.45
+CroCo ScanNet-50 86.60
+CroCo v2 5 datasets 51.55
+The results in Table˜7 show that Alligat0R consistently outperforms CroCo when trained on the same datasets, and even surpasses CroCo v2 which was pre-trained on 5 datasets.
+This demonstrates that our covisibility segmentation pre-training learns more generalizable features for dense correspondence tasks.
+A.8 Additional Qualitative Results
+We provide additional visualizations from our nuScenes and ScanNet datasets in Figures˜10 and 11, along with more predictions from Alligat0R and CroCo in Figure˜12.
+Figure 10: Covisibility annotation examples from Cub3 for nuScenes.
+For each image pair, we show the corresponding covisibility maps with color-coding for covisible, occluded, and outside FOV regions.
+Note how our annotation process handles varying degrees of overlap and challenging viewpoint changes.
+Let us highlight that some annotations, particularly the distinction between covisible and occluded pixels, may contain noise, especially for nuScenes, and we demonstrate in the experiments that Alligat0R is highly robust to this noise.
+Figure 11: Covisibility annotation examples from Cub3 for ScanNet.
+For each image pair, we show the corresponding covisibility maps with color-coding for covisible, occluded, and outside FOV regions.
+Note how our annotation process handles varying degrees of overlap and challenging viewpoint changes.
+ Reference    Target       Masked   CroCo recons.   Alligat0R segmentations
+Figure 12: Qualitative comparison.
+CroCo’s reconstructions are blurred in masked non-covisible regions, while Alligat0R often correctly identifies covisible, occluded, outside FOV regions across varying degrees of overlap and viewpoint changes.